The magnitudes of the gravitational field at distance r1 and r2 from the centre of a uniform sphere of radius R and mass m r F­1 and F2 respectively. Then :
 
(a) F1/F2 = r1/r2 if r1 < R and r2 < R
(b) F1/F2 = r22/r21 if r1 > R and r2 > R
(b) F1/F2 = r22 / r21 if r1 > R and r2 > R
(c) F1/F2 – r1 /r2 if r1 > R and r2 > R
(d) F2/F2 = r21 / r22 if r1 < R and r2 < R

Grade:6

1 Answers

Deepak Patra
9 years ago

(a,b,) For r > R, the gravitational field is F = GM / r2

? F1 = GM/r21 and F2 GM / r22 ? F1/F2 = r22 / r21

For r < R, the gravitational field F = GM/R3 x r

? F1 = GM/R3 x r1 and F2 = GM/R3 x r2

? F1/F2 = r1 / r2